# ProxyLM: Predicting Language Model Performance on Multilingual Tasks via Proxy Models

### v1

@@ Abstract
Performance prediction is a method to estimate the performance of multilingual language models (LMs), mitigating computational costs associated with model capacity and data for fine-tuning. Our paper introduces ProxyLM, a scalable framework for predicting LM performance using proxy models in multilingual tasks. These proxy models act as surrogates, approximating the performance of fine-tuned LMs on specific downstream natural language processing (NLP) tasks. By leveraging proxy models, ProxyLM significantly reduces computational overhead on task evaluations, achieving up to a 37.08 speedup compared to traditional methods, even with our smallest proxy models. Additionally, our methodology showcases adaptability to previously unseen languages in pre-trained LMs, outperforming the state-of-the-art performance by 1.89 as measured by root-mean-square-error (RMSE). This framework streamlines model selection, enabling efficient deployment and iterative LM enhancements without extensive computational resources.
@@ 1 Introduction
Given the limited academic computational resources for LM fine-tuning and inadequate LRL datasets, efficient methods in predicting model performance alleviate the dependency on extensive resources. While linear regression and gradient-boosting hold promise in performance prediction Birch et al. (2008a); Srinivasan et al. (2021); Xia et al. (2020a); Ye et al. (2021a); Schram et al. (2023); Khiu et al. (2024), existing solutions primarily focus on homogeneous data settings and prioritize high-resource languages using Transformer models Vaswani et al. (2017). Khiu et al. (2024) examine diverse datasets and LRLs but encounter limitations in the number of experiments, language diversity, and model scope, focusing solely on mBART Liu et al. (2020). Recent advancements in larger multilingual models, like NLLB Costa-jussà et al. (2022a) and M2M100 Fan et al. (2021), have significantly improved machine translation capabilities, exceeding those of mBART.
@@
We demonstrate the effectiveness and robustness of ProxyLM across dataset sources and languages on two estimated LM architectures. Our framework substantially outperforms all existing baselines in English-centric, many-to-many languages, and cross-dataset settings, including scenarios involving extremely LRLs that remain unseen by pre-trained LMs, surpassing the state-of-the-art performance measured with RMSE by .
@@
We also provide a time analysis comparing the fine-tuning duration of proxy models to direct model fine-tuning. Our results indicate that, with our smallest proxy models, we can achieve up to a speedup on task evaluation compared to the traditional approach, highlighting the efficiency of our approach.
@@ 2.1 ProxyLM
We present ProxyLM, a framework that leverages the past performance of other models, referred to as proxy models, as additional context for our regressor. Intuitively, proxy models can provide valuable insights that assist in predicting the performance of the estimated model . Formally, let be a set of proxy models. To integrate the information from these proxy models, we propose modifying Equation 2 as follows:
@@ 3 Experimental Setup
In this section, we describe the datasets and LMs used to obtain LMs’ performance records. These records were then used to train various regressor models under different experimental settings to investigate our approach to performance predictions. The details of the hyper-parameters for both the LMs and the regressors are provided in A.3.
@@ 3.1 Datasets
We evaluate our approach through two machine translation benchmarks: MT560 Gowda et al. (2021) and NusaTranslation Cahyawijaya et al. (2023). The MT560 dataset is English-centric, where English can serve as either the source or target language. We curated 20 datasets and selected 44 languages out of 500 for evaluation in MT560. In contrast, the NusaTranslation dataset comprises parallel texts in 12 Indonesian regional languages within a Many-to-Many Languages setting, allowing any language to act as the source or target. As many of these languages are absent in pre-trained multilingual models, we analyze 8 out of the 12 languages due to limited data in the remaining 4. The datasets encompass 50 languages across various domains such as economics, technology, and medicine. Detailed language insights are available in Tables 8 and 9 in A.1 for reference.
@@ 3.2 Models
We utilize four different transformer-based models: an encoder-decoder random initialized Transformers (100M) Vaswani et al. (2017), SMaLL-100 (330M) Mohammadshahi et al. (2022), M2M100 Fan et al. (2021), and NLLB Costa-jussà et al. (2022b). For M2M100 and NLLB, we use the models without any additional tuning (No FT) in a zero-shot fashion. Model details are provided in Appendix A.2. The evaluation is done using SentencePiece BLEU (spBLEU) Goyal et al. (2022), as it has been demonstrated to be a fair metric in multilingual settings, particularly in low-resource settings. For simplicity, the term “fine-tuning" will be used throughout this paper to refer to both the process of training from scratch (as in the case of the Transformer (100M) model) and the process of fine-tuning pre-trained LMs.
@@ 3.3 Experimental Settings
Each regressor is evaluated using RMSE as our performance metric and evaluated times with different seeds to obtain the mean and standard deviation of the performance results. We set our experiment settings as follows:
@@
Cross-Dataset: We train the regressor using performance records from the MT560 dataset and test it using records from the NusaTranslation dataset. We opt not to reverse this setup as the dataset exhibits no domain shift and contains fewer performance records.
@@ 4 Results and Analysis
In this section, we present the results of the performance predictions for ProxyLM and baselines over three settings: English-centric, Many-to-Many Languages, and Cross-Dataset, as described above. Further, we discuss the robustness, effectiveness, and efficiency of ProxyLM in the context of performance prediction.
@@ 4.1 English-centric Results
Table 1 shows the overall results on MT560. ProxyLM remarkably outperforms all existing baselines. We find that incorporating all proxy models (Ensemble) is the most effective for prediction, leading to a 2.29 averaged reduction in RMSE across all experimental settings compared to the best baseline. We observe that using the “No FT" estimated model to predict the performance of their fine-tuned models is surprisingly useful in all settings, especially for NLLB, where the model already has decent machine translation quality on LRLs. This observation is supported by our findings within the XGBoost model that the NLLB No FT feature has the highest importance score among all features, as shown in Figure 16. Further, using SMaLL-100 fine-tuned performance provides useful estimations for settings involving M2M100 as the estimated model. This may indicate that the performance of a model with similar architecture can be a good estimator for the performance of the larger estimated model. In other words, the choice of proxy model to help prediction matters. Feature importance analysis from the XGBoost model supports this, revealing that the SMaLL-100 fine-tuned feature has the highest importance score among all features, as shown in Figure 15.
Our analysis also indicates that XGBoost outperforms other regression models, across all evaluated settings. Both XGBoost and LGBM, which are gradient-boosting and tree-based learning methods, demonstrate superior performance metrics across all settings. Their robustness and efficiency as non-linear models are evident when compared to linear models, such as Poly2 and Poly3. Poly2 and Poly3 regressors, which employ second-degree and third-degree polynomial regression approaches respectively, tend to generate lower scores. This diminished performance is largely attributed to their limitations in capturing the nonlinear relationships inherent in the data, leading to suboptimal results. We further present the results by the language vitality on Table 2. The overall difference in RMSE between LRLs and medium-resource languages (MRLs) is relatively small, except when using SMaLL-100 as proxy models. An interesting observation here on SMaLL-100 in predicting the NLLB model is that the “No FT" model can predict “LRL" much better than the fine-tuned counterpart, and the fine-tuned model better predicts “MRL".
@@ 4.2 Many-to-Many Languages Results
Table 3 presents the performance of different models on the NusaTranslation dataset within the Many-to-Many languages setting. The results reveal that the Ensemble model achieves the lowest RMSE, with a 1.70 averaged reduction in RMSE across all experimental settings compared to the best baseline, indicating superior accuracy in performance predictions. An exception occurs in the random NLLB setting, where the model utilizing only NLPerf features outperforms the ensemble model, achieving the best performance. Note that no domain shift occurs within the dataset.
A comparative analysis shows that predicting the performance of M2M100 models in the random setting presents a greater challenge compared to predicting for the NLLB models. This discrepancy suggests that the complexity of performance prediction can vary substantially depending on the specific LM and the conditions under which it is evaluated. A particularly noteworthy finding is the effectiveness of using “No FT" models for estimating LM performance. The “No FT" models, which do not require any additional fine-tuning, demonstrate high accuracy in their performance predictions. This method offers substantial efficiency benefits, as it eliminates the need for extensive computational resources typically required for model training. In contrast, we find similar results between the LOLO setting for Many-to-Many languages and English-centric results, where ProxyLM using Ensemble remarkably outperforms all existing baselines. In addition, we find that using SMaLL-100 fine-tuned performance results in better predictions compared to those of the “No FT" estimated model.
@@ 4.3 Cross-Dataset Results
Table 4 illustrates model performance in the Cross-Dataset setup, showcasing the superior performance of ProxyLM with LGBM over XGBoost. The results highlight that ProxyLM with Ensemble significantly reduces RMSE compared to the best baseline by 2 and 1.69 for M2M100 and NLLB, respectively. This displays consistent performance across datasets and languages that were not encountered during the regressor’s training, including “unseen" languages for the pre-trained LMs. Moreover, the “No FT" models exhibit variability compared to other proxy models. The performance variation between M2M100 and NLLB may be attributed to the MT560 dataset solely containing "seen" languages for NLLB, lacking “unseen" languages examples for the regressor. This highlights the significance of incorporating “unseen" language instances in training for more dependable predictions.
@@ 4.4 Ablation Study
Figure 2 highlights the impact of features used in ProxyLM in the LOLO setting with XGBoost. Utilizing proxy models as features leads to a significant reduction in RMSE across all scenarios, showcasing their importance compared to other features. For the MT560 dataset, including language and dataset features alongside proxy models enhances performance. Dataset features alone show better improvement than language features alone, but the combination of both yields the best performance. On the other hand, for the NusaTranslation dataset, the benefits of incorporating dataset and language features are less pronounced, especially for the M2M100 model, and there may even be a performance dip for the NLLB model due to the dataset’s lack of domain shift.
@@ 4.5 Diminishing Returns with Increasing Training Set Size
In Figure 3, we examine the training of the XGBoost regressor using different numbers of MT560 past performance records as the training dataset. While the regressor’s performance shows enhancement with an expanding training size, the incremental benefits start diminishing once the training set surpasses about 400 past performance records. This observation implies that, across datasets, there exists a threshold where the advantages of incorporating additional past performance records begin to exhibit diminishing returns.
@@
| MT560 | NusaTranslation | ||||
@@
| Regressor Model | MT560 | NusaTranslation |
@@ 5 Related Work
The latter aims to predict the model performance by learning a correlation between input features and final evaluation metric. Birch et al. (2008a) identify strong predictive features such as the amount of reordering, the morphological complexity of the target language, and the historical relatedness of the two languages. Xia et al. (2020a) leverage extracted dataset features and typological database language representations. Ye et al. (2021a) introduce the use of confidence intervals and calibration with various regressor algorithms for reliable performance prediction. Schram et al. (2023) apply Bayesian matrix factorization for performance prediction on multilingual NLP tasks. In this work, we focus to explore the latter. Existing approaches have shown promise using linear regression and gradient-boosting trees Birch et al. (2008b); Xia et al. (2020b); Srinivasan et al. (2021); Ye et al. (2021b). These studies have considered data size, typological features, and language similarity as factors contributing to the model performance.
@@ 6 Conclusion
In this paper, we introduce ProxyLM, a novel framework designed to predict the performance of LMs by leveraging proxy models specifically for LRLs. By utilizing proxy models as substitutes to estimate the performance of the target model, we strategically employ smaller LMs and off-the-shelf models without additional fine-tuning. This framework is highly scalable to multiple proxy models and is task-agnostic, making it applicable to a wide range of downstream tasks. Our streamlined approach showcases substantial advancements in prediction accuracy compared to standard baselines and exhibits strong generalization capabilities across varied scenarios.
@@ Appendix A Experimental Details
| FLoRes Costa-jussà et al. (2022b) | afr, amh, arz, bak, bel, ceb, cym, dik, ewe, fao, | Multi-domain |
| guj, hau, hne, hye, ibo, ind, jav, kan, kat, kaz, | ||
| khm, kin, kir, lmo, ltz, mar, mri, mya, oci, pan, | ||
| plt, sin, sna, snd, som, ssw, tam, tat, tgl, tuk, | ||
| wol, xho, yor, zul | ||
@@ A.1 Languages Under Study
We list all the languages used in the training from the MT560 Gowda et al. (2021) and NusaTranslation Cahyawijaya et al. (2023) datasets in Table 8 and Table 9, respectively. The language code follows ∗ISO639-3 coding. All languages are also complemented by their †rarity taxonomy based on Joshi et al. (2021) into two vitality classes: 0-2→low resource language (LRL), and 3→mid resource language (MRL). We also provide information about whether the language was part of the pretrained M2M100 model dataset to highlight the model knowledge coverage.

### v2

@@ Abstract
Performance prediction is a method to estimate the performance of Language Models (LMs) on various Natural Language Processing (NLP) tasks, mitigating computational costs associated with model capacity and data for fine-tuning. Our paper introduces ProxyLM, a scalable framework for predicting LM performance using proxy models in multilingual tasks. These proxy models act as surrogates, approximating the performance of the LM of interest. By leveraging proxy models, ProxyLM significantly reduces computational overhead on task evaluations, achieving up to a 37.08 speedup compared to traditional methods, even with our smallest proxy models. Additionally, our methodology showcases adaptability to previously unseen languages in pre-trained LMs, outperforming the state-of-the-art performance by 1.89 as measured by root-mean-square error (RMSE). This framework streamlines model selection, enabling efficient deployment and iterative LM enhancements without extensive computational resources.
@@ 1 Introduction
Given the limited academic computational resources for LM fine-tuning and inadequate datasets for LRLs, performance prediction is an efficient method to alleviate the dependency on extensive resources by leveraging past performance records on an NLP task. While linear regression and gradient-boosting hold promise in performance prediction Birch et al. (2008a); Srinivasan et al. (2021); Xia et al. (2020a); Ye et al. (2021a); Schram et al. (2023); Khiu et al. (2024), existing solutions primarily focus on homogeneous data settings and prioritize high-resource languages using Transformer models Vaswani et al. (2017). Khiu et al. (2024) examine diverse datasets and LRLs but encounter limitations in the number of experiments, language diversity, and model scope, focusing solely on mBART Liu et al. (2020). Recent advancements in larger multilingual models, like NLLB Costa-jussà et al. (2022a) and M2M100 Fan et al. (2021), have significantly improved machine translation capabilities, exceeding those of mBART.
@@
We demonstrate the effectiveness and robustness of ProxyLM across 21 dataset sources and 50 languages on two estimated LM architectures. Our framework substantially outperforms all existing baselines in English-centric, many-to-many languages, and cross-dataset settings, including scenarios involving extremely LRLs that remain unseen by pre-trained LMs, surpassing the state-of-the-art performance measured with root-mean-square error (RMSE) by 1.89.
@@
We also provide a time analysis comparing the fine-tuning duration of proxy models to direct model fine-tuning. Our results indicate that, with our smallest proxy models, we can achieve up to a 37.08 speedup on task evaluation compared to the traditional approach, highlighting the efficiency of our approach.
@@ 2.1 ProxyLM
We present ProxyLM, a framework that leverages the past performance of other models, referred to as proxy models, as additional context for our regressor. Intuitively, proxy models can provide valuable insights that assist in predicting the performance of the estimated model . Formally, let be a set of proxy models. To integrate the information from these proxy models, we modify Equation 2 as follows:
@@ 3 Experimental Setup
In this section, we describe the datasets and LMs used to obtain LMs’ performance records. These records are then used to train various regressor models under different experimental settings to investigate our approach to performance predictions. The details of the hyper-parameters for both the LMs and the regressors are provided in A.3.
@@ 3.1 Datasets
We evaluate our approach using two types of datasets: English-centric and Many-to-Many Languages. The English-centric dataset involves English serving as either the source or target language. Our English-centric dataset comes from the MT560 Gowda et al. (2021) dataset, where we curated 19 datasets and selected 44 languages out of 500 for evaluation. Furthermore, we use the FLoRes-200 dataset Costa-jussà et al. (2022b) for additional validation and test sets. In contrast, the Many-to-Many Languages dataset allows any language to act as the source or target. We use the NusaTranslation dataset Cahyawijaya et al. (2023) as our Many-to-Many Languages dataset, which comprises parallel texts in 12 Indonesian regional languages. As many of these languages are absent in pre-trained multilingual models, we analyze 8 out of the 12 languages due to limited data in the remaining 4. Both datasets encompass 50 languages across various domains such as economics, technology, and medicine. Detailed language insights are available in Section A.1 in the Appendix.
@@ 3.2 Models
We utilize four different transformer-based models: an encoder-decoder random initialized Transformers (100M) Vaswani et al. (2017), SMaLL-100 (330M) Mohammadshahi et al. (2022), M2M100 Fan et al. (2021), and NLLB Costa-jussà et al. (2022b). For M2M100 and NLLB, we use the models without any additional tuning (No FT) in a zero-shot fashion. Model details are provided in Section A.2 in the Appendix. The evaluation is done using SentencePiece BLEU (spBLEU) Goyal et al. (2022), as it has been demonstrated to be a fair metric in multilingual settings, particularly in low-resource settings. For simplicity, the term “fine-tuning" will be used throughout this paper to refer to both the process of training from scratch (as in the case of the Transformer (100M) model) and the process of fine-tuning pre-trained LMs.
@@ 3.3 Experimental Settings
Each regressor is evaluated using RMSE as our performance metric and evaluated 5 times. We set our experiment settings as follows:
@@
Cross-Dataset: We train the regressor using performance records from the English-centric dataset and test it using records from the Many-to-Many Languages dataset. We opt not to reverse this setup as the dataset exhibits no domain shift and contains fewer performance records.
@@ 4 Results and Analysis
In this section, we present the results of the performance predictions for ProxyLM and baselines under the specified settings. Further, we discuss the robustness, effectiveness, and efficiency of ProxyLM in the context of performance prediction.
@@ 4.1 English-centric Results
Table 1 shows the overall results on the English-centric dataset. ProxyLM remarkably outperforms all existing baselines. We find that incorporating all proxy models (Ensemble) is the most effective for prediction, leading to a 2.29 averaged reduction in RMSE across all experimental settings compared to the best baseline. We observe that using the No FT estimated model to predict the performance of their fine-tuned models is surprisingly useful in all settings, especially for NLLB, where the model already has decent machine translation quality on LRLs. This observation is supported by our findings within the XGBoost model that the NLLB No FT feature has the highest importance score among all features, as shown in Figure 16 in the Appendix. Further, using SMaLL-100 fine-tuned performance provides useful estimations for settings involving M2M100 as the estimated model. This may indicate that the performance of a model with similar architecture can be a good estimator for the performance of the larger estimated model. In other words, the choice of proxy model to help prediction matters. Feature importance analysis from the XGBoost model supports this, revealing that the SMaLL-100 fine-tuned feature has the highest importance score among all features, as shown in Figure 15 in the Appendix.
Our analysis also indicates that XGBoost outperforms other regression models, across all evaluated settings. Both XGBoost and LGBM, which are gradient-boosting and tree-based learning methods, demonstrate superior performance metrics across all settings. Their robustness and efficiency as non-linear models are evident when compared to linear models, such as Poly2 and Poly3. Poly2 and Poly3 regressors, which employ second-degree and third-degree polynomial regression approaches respectively, tend to generate lower scores. This diminished performance is largely attributed to their limitations in capturing the nonlinear relationships inherent in the data, leading to suboptimal results. We further present the results by the language vitality in Table 2. The overall difference in RMSE between LRLs and medium-resource languages (MRLs) is relatively small, except when using SMaLL-100 as proxy models. An interesting observation here on SMaLL-100 in predicting the NLLB model is that the No FT model can predict LRLs much better than the fine-tuned counterpart, and the fine-tuned model better predicts MRLs.
@@ 4.2 Many-to-Many Languages Results
Table 3 presents the performance of different models on the Many-to-Many Languages dataset. The results reveal that the Ensemble model achieves the lowest RMSE, with a 1.70 averaged reduction in RMSE across all experimental settings compared to the best baseline, indicating superior accuracy in performance predictions. An exception occurs in the random NLLB setting, where the model utilizing only NLPerf features outperforms the ensemble model, achieving the best performance. Note that no domain shift occurs within the dataset.
A comparative analysis shows that predicting the performance of the M2M100 model in the random setting presents a greater challenge compared to predicting the NLLB model. This discrepancy suggests that the complexity of performance prediction can vary substantially depending on the specific LM and the conditions under which it is evaluated. A particularly noteworthy finding is the effectiveness of using No FT models for estimating LM performance. The No FT models, which do not require any additional fine-tuning, demonstrate high accuracy in their performance predictions. This method offers substantial efficiency benefits, as it eliminates the need for extensive computational resources typically required for model training. In contrast, we find similar results between the LOLO setting for Many-to-Many Languages and English-centric results, where ProxyLM using Ensemble remarkably outperforms all existing baselines. In addition, we find that using SMaLL-100 fine-tuned performance results in better predictions compared to those of the No FT estimated model.
@@ 4.3 Cross-Dataset Results
Table 4 illustrates model performance in the Cross-Dataset setup, showcasing the superior performance of ProxyLM with LGBM. We use LGBM instead of XGBoost because LGBM performs better. The results highlight that ProxyLM with Ensemble significantly reduces RMSE compared to the best baseline by 2 and 1.69 for M2M100 and NLLB, respectively. This displays consistent performance across datasets and languages that are not encountered during the regressor’s training, including unseen languages for the pre-trained LMs. However, the No FT models exhibit inconsistency compared to other proxy models. We also find varied performance between M2M100 and NLLB, which may be attributed to the English-centric dataset containing only seen languages for NLLB and lacking unseen language examples for the regressor. This highlights the significance of incorporating unseen language instances in training for more robust predictions.
@@ 4.4 Ablation Study
Figure 2 highlights the impact of features used in ProxyLM in the LOLO setting with XGBoost. Utilizing proxy models as features leads to a significant reduction in RMSE across all scenarios, showcasing their importance compared to other features. For the English-centric dataset, including language and dataset features alongside proxy models enhances performance. Dataset features alone show better improvement than language features alone, but the combination of both yields the best performance. On the other hand, for the Many-to-Many Languages dataset, the benefits of incorporating dataset and language features are less pronounced, especially for the M2M100 model, and there may even be a performance dip for the NLLB model due to the dataset’s lack of domain shift.
@@ 4.5 Diminishing Returns with Increasing Training Set Size
In Figure 3, we examine the training of the XGBoost regressor using different numbers of English-centric past performance records as the training dataset. While the regressor’s performance shows enhancement with an expanding training size, the incremental benefits start diminishing once the training set surpasses about 400 past performance records. This observation implies that, across datasets, there exists a threshold where the advantages of incorporating additional past performance records begin to exhibit diminishing returns.
@@
| English-centric | Many-to-Many Langs. | ||||
@@
| Regressors | English-centric | Many-to-Many Langs. |
@@ 5 Related Work
The latter aims to predict the model performance by learning a correlation between input features and the final evaluation metric. Birch et al. (2008a) identify strong predictive features such as the amount of reordering, the morphological complexity of the target language, and the historical relatedness of the two languages. Xia et al. (2020a) leverage extracted dataset features and typological database language representations. Ye et al. (2021a) introduce the use of confidence intervals and calibration with various regressor algorithms for reliable performance prediction. Schram et al. (2023) apply Bayesian matrix factorization for performance prediction on multilingual NLP tasks. In this work, we focus on exploring the latter. Existing approaches have shown promise using linear regression and gradient-boosting trees Birch et al. (2008b); Xia et al. (2020b); Srinivasan et al. (2021); Ye et al. (2021b). These studies have considered data size, typological features, and language similarity as factors contributing to the model performance.
@@ 6 Conclusion
In this paper, we introduce ProxyLM, a novel framework designed to predict the performance of LMs by leveraging proxy models including for LRLs. By utilizing proxy models as substitutes to estimate the performance of the target model, we strategically employ smaller LMs and off-the-shelf models without additional fine-tuning. This framework is highly scalable to multiple proxy models and is task-agnostic, making it applicable to a wide range of downstream NLP tasks. Our streamlined approach showcases substantial advancements in prediction accuracy compared to standard baselines and exhibits strong generalization capabilities across varied scenarios.
@@ Appendix A Experimental Details
| English-centric Dataset | ||
| FLoRes-200 Costa-jussà et al. (2022b) | afr, amh, arz, bak, bel, ceb, cym, dik, ewe, fao, | Multi-domain |
| guj, hau, hne, hye, ibo, ind, jav, kan, kat, kaz, | ||
| khm, kin, kir, lmo, ltz, mar, mri, mya, oci, pan, | ||
| plt, sin, sna, snd, som, ssw, tam, tat, tgl, tuk, | ||
| wol, xho, yor, zul | ||
@@
| Many-to-Many Languages Dataset | ||
@@ A.1 Languages Under Study
We list all the languages used in the training from the MT560 Gowda et al. (2021) and NusaTranslation Cahyawijaya et al. (2023) datasets in Table 8 and Table 9, respectively. The language code follows ∗ISO639-3 coding. All languages are also complemented by their †rarity taxonomy based on Joshi et al. (2021) into two vitality classes: 0-2→low resource language (LRL), and 3→mid resource language (MRL). We also provide information about whether the language was part of the pre-trained M2M100 model dataset to highlight the model knowledge coverage.